# Generative Tools for Graphical Assets: Empirical Guidelines based on Game Designers’ and Developers’ Preferences

### v1

@@ Abstract
[…]
@@ 1 Introduction
It is apparent that different generative methods and techniques are designed with different use cases in mind, each attempting to solve a problem or streamline a process within a larger creative pipeline. Practitioners use many tools during game design and production, from editor tools for game engines, to asset editing software such as Photoshop (Adobe Inc., ) and Blender (Blender Foundation, 2024), or integrated development environments such as Visual Studio (Microsoft Corporation, ). These tools each have their place within the workflows and development pipelines of game creators. While these workflows look different from company to company or from one individual to another, there exists three ubiquitous stages of development: prototyping/design, production and testing (Ramadan and Widyani, 2013). All graphical assets are formulated and produced during the prototyping, design and production stages. Within this, there are many ways that PCG can help, from early inspiration and creating rough placeholders to creating fully fledged assets or remixing existing assets for variety. Each of these use cases, naturally, will have different requirements for the complexity of input, method of interaction and quality of output. However, no research has yet explored the purposes for which designers and developers prefer to use PCG tools.
The user study of Walton et al. (Walton et al., 2021), examines user opinions of a MI-PCG level creation tool using a mixed-methods approach. The findings suggest that the MI-PCG method stimulates user creativity, and helps to inspire new ideas. They discover that the inclusion of qualitative data from participants is important for gaining the full picture, as it provides important context to the quantitative results.
@@
While there is an extensive body of literature pertaining to MI-PCG there are limited case studies which examine their use and fit in design and development pipelines (Lai et al., 2020). While it is clear that MI-PCG has the potential to improve game development asset pipelines, the opinions and preferences of game designers and developers remain, in general, unheard. Though recent work examines usage applications for generative AI systems such as ChatGPT (Yan et al., 2024) among general user bases, the specific needs and requirements within the video game industry have not yet been addressed. Individual MI-PCG systems may be validated with user feedback (Li et al., 2020; Shen et al., 2020; Shen and Chen, 2021), but feedback is limited to specific purposes of use, and are not contextualised within the larger scope of the development pipeline. The goal of this study is to obtain insights into where and how graphical asset PCG is most useful to game designers and developers.
Section 2 will expand on the motivation for this study, followed by section 3 in which the seven pipeline applications will be identified. Section 4 will then introduce the research questions. Section 5 will describe the study procedure, followed by sections 6 and 7 which will explain the mock-ups that were tested and the experiment results respectively. Section 8 will discuss the findings and present guidelines for creators of graphical asset generative tools, and further research.
2 Motivation
@@
We propose a list of potential usages for graphical asset generation within design and development pipelines, shown in figure 2, along with key characteristics. These characteristics include whether the implementation is expected to be offline or online, whether the outputs are expected to be high quality or low quality, and if the generator supports an existing development process, or replaces it. This designation will map the potential usages to the choice of approach within the GaGeTx framework.
@@
3 Applications of GAGs in game design and development pipelines
The creation of graphical assets occurs during the prototyping, design and production phases of the game design and development process. Graphical asset generation may be utilised throughout each of these phases to streamline or augment the creation of games. Through observing existing generative methods, tools and pipelines in the literature, seven potential usages for graphical asset generators have been identified:
@@ 4 Research Questions
The aim of this study is two-fold: a) to establish if such generative systems would be found useful by game designers and developers, and what their requirements are for such a system, and b) to discover where in the the pipeline such tools would best fit, and how the chosen generative method may line up with this. As a result, this will provide expand and provide nuance to the GAGeTx framework.
@@ 5 Procedure
To obtain preferences of game developers and designers with regard to generative tools for graphical assets, we developed three mock-up interfaces, designed to cover broad forms of interaction. Each participant was asked to complete a form consisting of demographic questions, followed by a section for each mock-up. In each section the mock-up is introduced and instructions for installing and using the mock-up are provided. The participant is asked to try the mock-up for as long as they like, then complete a set of preference questions. After all three sections are completed, the participant is provided with an opportunity to book a semi-structured interview to provide qualitative feedback. In analysis, the questionnaire responses are compared between the three mock-ups, and overall. The following subsections will provide more detail on the participants, mock-ups, questionnaires and interviews.
@@ 5.1 Participants
Participants were recruited via a convenience sampling approach, in which members of the game development communities: LUUG and BCS Animation and Games specialist group, were contacted via group emails through the respective community administrators. Participants were required to be over the age of 18, with professional experience in game design or development. Participants were also required to have Unity engine installed on their personal machines, and experience with the software was preferred. Of the group email respondents, participants were recruited based on the aforementioned criteria. Participation was not incentivised on the basis of financial gain, and was thus voluntary.
@@ 6 Mock-ups
The three mock-ups represent stand-alone wizard, editor window integrated and editor inspector integrated implementations of this system. Figure 3 shows an example usage process implemented in each of the three mock-ups. This process involves the configuration of a generator that takes text as input, and outputs bitmap images.
@@
step 1
step 2
step 3
step 4
step 5
M1
step 1
step 2 & 3
step 4
step 5
M2
step 1
step 2 & 3
step 4
step 5 part 1
step 5 part 2
M3
@@ 7.5 Findings
For H1, given answers to RQ5, we find that users prefer an integrated solution over a stand-alone implementation, with a mean value of .78 ($\sigma$ = .428), where 1 represents an ‘integrated solution’. This is confirmed by RQ1, in which the integrated solutions scored higher. With regards to H2, results of RQ1 and RQ6.1 suggest no significant preference between window and inspector integrated interfaces. H4 is confirmed with a preference for volume over quality across the board, with a mean value of .17 ($\sigma$ = .383), where 0 represents ‘volume over quality’. Regarding H5 (RQ4), a mean selection of 1.67 ($\sigma$ = .686) shows a preference for a maximum generation time between 1 and 10 minutes for a single asset, rejecting this hypothesis. For H6, a mean score of 3.5 ($\sigma$ = .985) on a 5-point Likert scale, RQ6, suggests a moderate preference for the ability to configure or modify the tool. Answers to RQ7 present an overall mean SUS score of 78.75 ($\sigma$ = 13.429), therefore confirming H7. The individual mean SUS scores for M1, M2 and M3 were 80.00 ($\sigma$ = 12.649), 78.75 ($\sigma$ = 10.694), and 77.50 ($\sigma$ = 18.303) respectively. Following the adjective rating system provided by Bangor et al. (Bangor et al., 2009), the mean overall score can be classed as "Good" (78.75). Section 8 will discuss these findings and provide recommendations, as well as suggestions for future research.
8 Discussion of Results: Guidelines and Recommendations for Generative Tools
@@
Figure 5: Guidelines resulting from research findings.
@@
Overall, results are in alignment with the findings of Walton et al. (Walton et al., 2021) and Kasurinen et al. (Kasurinen et al., 2013). Users were interested in using the tool for inspiration and prototyping, and favored integration, and thus compatibility with their chosen game engine. Additionally, when observing the characteristics of the three preferred usages according to the characteristics shown in figure 2, in general, users prefer generators for work in an offline context, with an expectation for lower quality and to augment, rather than replace an existing process in their pipeline.
@@
Figure 4 presents the flow of graphical asset data in both stand-alone and integrated scenarios. As shown, creative tools can either be integrated within the main development environment (game engine) or implemented externally as stand-alone programs. Users interact with these tools via user-interfaces, and in the case of integrated tools, the implementation itself can rely on the native features of the environment through application programming interfaces (API). Stand-alone external generative tools, as with other creation tools such as Photoshop (Adobe Inc., ), Blender (Blender Foundation, 2024) and Visual Studio (Microsoft Corporation, ) instead rely on outputting artefacts in standardised formats that most game engines support, such as FBX, OBJ, PNG and JPEG. Integrated tools may implement GAG technique interactions and processes using game engine UI APIs and native engine features via back end APIs respectively.
@@
Figure 5 presents guidelines resulting from the findings of this user study, spanning both the user intent as well as the implementation and integration of graphical asset generators in design/development pipelines. Table 9 presents the mapping of these guidelines to the research questions.
Table 9: Mapping of guidelines to research questions.
Recommendations Research Questions
Facilitate tool configurability RQ6
Cater to early pipeline usage RQ2, RQ3, RQ4
Augment a stage of design and development RQ2
Match the design language of the environment RQ8
Use common and expected data formats RQ8, RQ9
Develop a suitable interface for the underlying interactions RQ1, RQ7
Integrate with application programming interfaces (APIs) RQ5, RQ8
@@
The first consideration is user intent with regard to the usage application, what technique interaction and process is used and what the desired outcome is. When considering user intent, based on the findings, there are three main considerations:
@@
  • Cater to early pipeline usage. Users favour early pipeline generation for purposes of generating inspiration, exploring ideas, and creating placeholders, as presented in figure 5. Furthermore, the system should aim for variety and volume of outputs rather than quality. This is because users prefer not to use the results in a finished product directly, but would rather use them as ideas to build on and refine.
@@
  • Augment a stage of design and development. The preferred use cases are all supportive of a human-driven design and development process, and do not directly replace the final creation of assets. Rather, they give designers and developers ideas to expand on, break creative block or help to streamline more mundane or inconsequential tasks. This entails the majority of applications, excluding player made content as shown in figure 5.
@@
  • Facilitate tool configurability. Game design and development is a creative endeavor and thus, new, unique forms of asset are part and parcel. A graphical asset generator that has a pre-defined style or rigid way of working becomes a tool with limited applicability. Configurability should be considered, regardless of the chosen application.
@@
Once intent is established, and the GAG technique and approach is determined, the tool must be implemented and integrated within the game design and development pipeline. Here there are four main considerations:
@@
  • Match the design language of the environment. The less the user has to learn on top of what they are accustomed to within their pipeline, the easier it is to incorporate the tool. If the interface of the tool provides functionality that other native features of the environment have, such as the ability to undo and redo changes, frustration or confusion can be avoided. This can be achieved through appropriate use of front-end UI APIs within the host software, as illustrated in figure 4.
@@
  • Use common and expected data formats. To smoothly integrate the tool, whether it is integrated or stand-alone, the output artefacts must be in a format that the development environment can utilise. Typically these are ubiquitous data formats, such as OBJ or common proprietary formats that are supported due to popularity, such as Autodesk FBX. This is particularly important in the transfer of artefacts from and between external generative tools, creation tools and the users development environment of choice, as shown in figure 4.
@@
  • Develop a suitable interface for the underlying interactions. Design an interface to best interact with the features of the tool, assuming that it provides good usability, there is no preference between step by step options, graph view windows and inspector editors.
@@
  • Integrate with application programming interfaces (APIs). The tool should be integrated as a part of an existing popular engine or editor, or at the very least should have full compatibility with said application. Technique interactions and processes must interact with each other while appropriately utilising the host game engine’s front-end and back-end APIs, as shown in figure 4.

### v2

# Heuristics for AI-driven Graphical Asset Generation Tools in Game Design and Development Pipelines: A User-Centred Approach

@@ Abstract
Graphical assets play an important role in the design and development of games. There is potential in the use of AI-driven generative tools, to aid in creating graphical assets, thus improving game design and development pipelines. However, there is little research to address how the generative methods can fit into the wider pipeline. There also no guidelines or heuristics for creating such tools. To address this gap we conducted a user study with 16 game designers and developers to examine their behaviour and interaction with generative tools for graphical assets. The findings highlight that early design stage is preferred by all participants. Designers and developers are inclined to use such tools for creating large amounts of variations at the cost of quality as they can improve the quality of the artefacts once they generate a suitable asset. The results also strongly raised the need for better integration of such tools in existing design and development environments and the need for the outputs to be in common data formats, to be manipulatable and smoothly integrate into existing environments. The study also highlights the requirement for further emphasis on the needs of the users to incorporate these tools effectively in existing pipelines. Informed by these results, we provide a set of heuristics for creating tools that meet the expectations and needs of game designers and developers.
@@ 1 Introduction
It is apparent that different generative methods and techniques are designed with different use cases in mind, each attempting to solve a problem or streamline a process within a larger creative pipeline. Practitioners use many tools during game design and production, from editor tools for game engines, to asset editing software such as Photoshop and Blender, or integrated development environments such as Visual Studio. These tools each have their place within the workflows and development pipelines of game creators. While these workflows look different from company to company or from one individual to another, there exists three ubiquitous stages of development: prototyping/design, production and testing (Ramadan and Widyani, 2013). All graphical assets are formulated and produced during the prototyping, design and production stages. Within this, there are many ways that PCG can help, from early inspiration and creating rough placeholders to creating fully fledged assets or remixing existing assets for variety. Each of these use cases, naturally, will have different requirements for the complexity of input, method of interaction and quality of output. However, no research has yet explored the behaviour and interaction patterns of the users, i.e. the purposes for which designers and developers use PCG tools and how they fit in their pipelines.
While there is an extensive body of literature pertaining to MI-PCG there are limited case studies which examine their use and fit in design and development pipelines (Lai et al., 2020). While it is clear that MI-PCG has the potential to improve game development asset pipelines, the behaviour, opinions and preferences of game designers and developers remain, in general, unheard. Though recent work examines usage applications for generative AI systems such as ChatGPT (Yan et al., 2024) among general user bases, the specific needs and requirements within the video game industry and research have not yet been addressed. Individual MI-PCG systems may be validated with user feedback (Li et al., 2020; Shen et al., 2020; Shen and Chen, 2021), but feedback is limited to specific purposes of use, and are not contextualised within the larger scope of the development pipeline. The goal of this study is to obtain insights into where and how graphical asset PCG is most useful to game designers and developers.
The following section, explores the current state of research in this area. Section 3 will expand on the motivation for this study and identify seven pipeline applications. Section 4 will then introduce the research questions. Section 5 will describe the study procedure, followed by sections 6 and 7 which will explain the mock-ups that were tested and the experiment results respectively. Sections 8 and 9 will discuss the findings and present our heuristics for creators of graphical asset generative tools, and further research.
2 Review of Current Research
The process of creating games is complex, requiring design decisions at every level from the overall concept and gameplay down to the characters and environments. Designers and developers make use of increasingly capable creative tools to streamline this task, which they often tailor to their own needs (Newell et al., 2021; Seidel et al., 2019).
The vast amount of content required in video games requires an equally vast amount of time and resources to create. Generative approaches can be applied to reduce this work load (Seidel et al., 2019).
It is important for designers and developers to have control over the content that is created, thus generative tools typically involve mixed-initiative (MI) approaches. The user study of Walton et al. (Walton et al., 2021), examines user opinions of a PCG level creation tool using a mixed-methods approach. The findings suggest that the MI-PCG method stimulates user creativity, and helps to inspire new ideas. They discover that the inclusion of qualitative data from participants is important for gaining the full picture, as it provides important context to the quantitative results.
@@ 2 Review of Current Research
Sketchar (Ling et al., 2024) is a mixed-initiative generative tool encompassing a full character design pipeline. In this method, users iteratively build on each aspect of a design, including character profiles, concept art, relationships with other characters and dialogue between them, using large language models (LLMs) such as ChatGPT and image generators such as DALL-E. In user testing, designers identify the benefits of this method in elaborating on existing, early-stage ideas as well as visually communicating their design ideas to artists.
Colado et al. (Colado et al., 2023), examine mixed-initiative generative approaches to serious game design. Much like Sketchar (Ling et al., 2024), this method implemented an iterative text-to-image generation workflow, this time for character concepts as well as scenarios. Results from initial testing suggests that this method can save time in development and aid smaller teams with less resources.
It is clear that generative methods have clear applications in graphical asset pipelines. According to Unity’s game report in 2024, 62% of developers that have adopted AI tools use them for asset generation (Unity Technologies, 2024). However, the impact of generative, mixed-initiative tools for graphical game assets is yet to be examined.
3 Motivation & Current Development Pipeline
@@
Through the analysis of existing game development pipelines, Ramadan and Widyani, (Ramadan and Widyani, 2013) introduced a comprehensive game development life cycle (GDLC). In this framework, practitioners begin with a rough concept of the game then begin a design and development loop starting with pre-production, in which more detailed concepts are established, then followed by production, in which the concepts are implemented, and testing in which these implementations are appraised. Insights from testing then feed into further pre-production and the cycle continues, iterating on the design until the practitioners are satisfied and they move on to refinement in the beta phase, then finally release the product.
To build on this, we present an expanded version of the GDLC, incorporating the graphical asset creation pipeline as it relates to general development, shown in figure 3. Here we identify points in the pipeline at which practitioners utilise creative tools, through observing existing generative methods, tools and pipelines in the literature. These include building inspiration in the interim between initiation and pre-production; generating placeholder assets, and exploring designs during pre-production and early production; implementing these designs as full (Core) assets in production; creating variations of these assets in the refinement stage; and supporting player/user made content.
@@
Figure 3: Current game design and development pipeline, adapted from (Ramadan and Widyani, 2013).
@@ 4 Research Questions
The aim of this study is two-fold: a) to establish if such generative tools would be found useful by game designers and developers, and what their requirements are for interacting with such tools, and b) to discover where in the pipeline such tools would best fit, and how the chosen generative method may line up with this. As a result, this will provide expand and provide nuance to the GAGeTx framework.
@@ 5 Procedure
To obtain preferences of game developers and designers with regard to generative tools for graphical assets, we developed three mock-up interfaces, designed to cover broad forms of interaction. Each participant was asked to complete a form consisting of demographic questions, followed by a section for each mock-up. In each section the mock-up is introduced and instructions for installing and using the mock-up are provided. The participant is asked to try the mock-up for as long as they like, then complete a set of preference questions. After all three sections are completed, the participant is provided with an opportunity to book a semi-structured interview to provide qualitative feedback. In analysis, the questionnaire responses are compared between the three mock-ups, and overall. The study was approved by University’s Research Ethics Committee (37113-LR-Jun/2022-39654-1). The following subsections will provide more detail on the participants, mock-ups, questionnaires and interviews.
@@ 5.1 Participants
Participants were recruited via a convenience sampling approach, in which members of the game development communities: LUUG and BCS Animation and Games specialist group, were contacted via group emails through the respective community administrators. Participants were required to be over the age of 18, with professional experience in game design or development. Participants were also required to have Unity engine installed on their personal machines, and experience with the software was preferred. Of the group email respondents, participants were recruited based on the aforementioned criteria, and signed a consent form before taking part the study. Participation was not incentivised on the basis of financial gain, and was thus voluntary.
@@ 6 Mock-ups
The three mock-up interfaces represent stand-alone wizard, editor window integrated and editor inspector integrated implementations of this system. Figure 4 shows an example usage process implemented in each of the three mock-ups. This process involves the configuration of a generator that takes text as input, and outputs bitmap images.
@@
Creating a new graphical asset generator
Defining target asset type
Defining target format
Assigning the technique
Assigning input and output sources
Stand-alone wizard interface (M1)
Creating a new empty graphical asset generator graph
Defining target asset type and format
Assigning the technique
Assigning input and output sources
Integrated in an editor window interface (M2)
Creating a new graphical asset generator
Defining target asset type and format
Assigning the technique
Assigning input sources
Assigning output sources
Integrated in an editor inspector view interface (M3)
@@ 7.5 Findings
For H1, given answers to RQ5, we find that users prefer an integrated solution over a stand-alone implementation, with a mean value of .78 ($\sigma$ = .428), where 1 represents an ‘integrated solution’. This is confirmed by RQ1, in which the integrated solutions scored higher. With regards to H2, results of RQ1 and RQ6.1 suggest no significant preference between window and inspector integrated interfaces. H4 is confirmed with a preference for volume over quality across the board, with a mean value of .17 ($\sigma$ = .383), where 0 represents ‘volume over quality’. Regarding H5 (RQ4), a mean selection of 1.67 ($\sigma$ = .686) shows a preference for a maximum generation time between 1 and 10 minutes for a single asset, rejecting this hypothesis. For H6, a mean score of 3.5 ($\sigma$ = .985) on a 5-point Likert scale, RQ6, suggests a moderate preference for the ability to configure or modify the tool. Answers to RQ7 present an overall mean SUS score of 78.75 ($\sigma$ = 13.429), therefore confirming H7. The individual mean SUS scores for M1, M2 and M3 were 80.00 ($\sigma$ = 12.649), 78.75 ($\sigma$ = 10.694), and 77.50 ($\sigma$ = 18.303) respectively. Following the adjective rating system provided by Bangor et al. (Bangor et al., 2009), the mean overall score can be classed as "Good" (78.75). The next sections will discuss these findings and propose a set of heuristics, as well as suggestions for future research.
Figure 5: The game design and development pipeline based on the GDLC of Ramadan and Widyani (Ramadan and Widyani, 2013) and expanded to include graphical asset findings.
8 Discussion and Comparison of Results
@@
Users were interested in using the tool for inspiration and prototyping, and favoured integration, and thus compatibility with their chosen game engine. Additionally, when observing the characteristics of the three preferred usages according to the characteristics shown in figure 2, in general, users prefer generators for work in an offline context, with an expectation for lower quality and to augment, rather than replace an existing process in their pipeline.
Given the findings regarding pipeline preferences, we refine the graphical asset GDLC to include the offline generative tool applications, shown in figure 5. Here, the pipeline applications creating assets from existing designs and creating assets from scratch both apply under the Assets (core) step and have thus been combined. We also show the human-computer interaction for each of these steps with early pipeline tasks: inspiration, placeholders and explore, employing an iterative shared initiative approach, where both human and AI provide equal input; and late pipeline tasks: Assets (core) and Assets (vary), taking a human-driven approach, in which designs are already well defined and the goal is to implement them in a more controlled and selective manner. This addition formally situates the well-corroborated preference for applying generative tools in the early stages of the pipeline.
Our results align with and expand upon existing research findings regarding game design and development tools (Walton et al., 2021; Kasurinen et al., 2013; Ling et al., 2024; Colado et al., 2023; Togelius et al., 2011; Lai et al., 2020), addressing the gaps in previous research. Namely, previous work has suggested that designers and developers value variety in outputs (Walton et al., 2021), which corroborates with the usage of generative methods as explorative tools (Ling et al., 2024; Colado et al., 2023; Kasurinen et al., 2013). Our results not only reflect this need and preference, but identify the exact use cases designers and developers target within the pipeline, providing further context to this requirement and allowing tools to cater to those specific needs.
Although previous research demonstrates an interest in using these tools in pipelines (Ling et al., 2024; Colado et al., 2023), they do not clarify how and where in the pipeline GAG tools would be suitable. Our findings show that they are much more preferred in the early stages. Furthermore, we expand on research that suggests designers and developers prefer well integrated configurable tools (Kasurinen et al., 2013) through direct quantitative results and interview feedback emphasising compatibility with their tools. Our results show that the style of interface only matters as far as it suits the interactions needed for the technique in use. To integrate these interactions into existing software the interface must be achieved through APIs, which also inform the design language of the tool.
Previous works focus on high-level tool implications (Togelius et al., 2011; Walton et al., 2021), but do not consider the practical aspects that can make or break the applicability of a tool. On the contrary, Lai et al. (Lai et al., 2020) identify a need to respect existing work processes. We elaborate on this by pinpointing the use of standardised data formats as an important requirement for generative tools in practice, as well as the need for design features that align with the host environment and thus meet user expectations.
While existing work acknowledges that different balances of user-machine initiative can be useful in different situations (Liapis et al., 2016; Lai et al., 2020), research has not yet ascertained how this balance maps onto the stages of a design and development pipeline. Through our findings we identify the appropriate flow of initiative at different stages of a graphical asset pipeline as presented in figure 5. At earlier stages of the pipeline less user control is required, thus the tool can have a larger share of the initiative. Once development shifts into a production phase, the design choices have already been made, and thus the user takes a more prescriptive role over the tool. This helps to inform the design and integration of GAG tools, based on their intended stage in the pipeline.
@@
The flow of graphical asset data in both stand-alone and integrated scenarios is presented in figure 6. As shown, creative tools can either be integrated within the main development environment (game engine) or implemented externally as stand-alone programs. Users interact with these tools via user-interfaces, and in the case of integrated tools, the implementation itself can rely on the native features of the environment through application programming interfaces (API). Stand-alone external generative tools, as with other creation tools such as Photoshop (Adobe Inc., 2005), Blender (Blender Foundation, 2024) and Visual Studio (Microsoft Corporation, 2022) instead rely on outputting artefacts in standardised formats that most game engines support, such as FBX, OBJ, PNG and JPEG. Integrated tools may implement GAG technique interactions and processes using game engine UI APIs and native engine features via back end APIs respectively.
9 Heuristics for Generative Tools
@@
To improve the applicability of our findings to the design and integration of GAG tools, we have derived a set of actionable heuristics, shown in figure 7, spanning both the user intent as well as the implementation and integration of GAGs in design/development pipelines. Table 9 presents the mapping of these heuristics to the research questions.
Figure 7: Mapping the user considerations, intent and implementation/integration to the heuristics.
Table 9: Mapping of heuristics to research questions.
Heuristics Research Question(s)
H-CON Facilitate tool configurability RQ6
H-PIP Cater to early pipeline usage RQ2, RQ3, RQ4
H-SYN Synergise generative and human-driven design and development RQ2
H-LAN Match the design language of the environment RQ8
H-DAT Use common and expected data formats RQ8, RQ9
H-INT Develop interfaces to suit the underlying interactions RQ1, RQ7
H-API Integrate with application programming interfaces (APIs) RQ5, RQ8
@@
The first consideration is user intent with regard to the usage application, what technique interaction and process is used and what the desired outcome is. When considering user intent, based on the findings, there are three heuristics:
@@
  • H-PIP: Cater to early pipeline usage. Users favour early pipeline generation for purposes of generating inspiration, exploring ideas, and creating placeholders, as presented in figure 7. Furthermore, the system should aim for variety and volume of outputs rather than quality. This is because users prefer not to use the results in a finished product directly, but would rather use them as ideas to build on and refine.
@@
  • H-SYN: Synergise generative and human-driven design and development. The preferred use cases are all supportive of a human-driven design and development process, and do not directly replace the final creation of assets. Rather, they give designers and developers ideas to expand on, break creative block or help to streamline more mundane or inconsequential tasks. This entails the majority of applications, excluding player made content as shown in figure 7.
@@
  • H-CON: Facilitate tool configurability. Game design and development is a creative endeavor and thus, new, unique forms of assets are part and parcel. A graphical asset generator that has a pre-defined style or rigid way of working becomes a tool with limited applicability. Configurability should be considered, regardless of the chosen application.
@@
Once intent is established, and the GAG technique and approach is determined, the tool must be implemented and integrated within the game design and development pipeline. Here, there are four heuristics:
@@
  • H-LAN: Match the design language of the environment. The less the user has to learn on top of what they are accustomed to within their pipeline, the easier it is to incorporate the tool. If the interface of the tool provides functionality that other native features of the environment have, such as the ability to undo and redo changes, frustration or confusion can be avoided. This can be achieved through appropriate use of front-end UI APIs within the host software, as illustrated in figure 6.
@@
  • H-DAT: Use common and expected data formats. To smoothly integrate the tool, whether it is integrated or stand-alone, the output artefacts must be in a format that the development environment can utilise. Typically these are ubiquitous data formats, such as OBJ or common proprietary formats that are supported due to popularity, such as Autodesk FBX. This is particularly important in the transfer of artefacts from and between external generative tools, creation tools and the users development environment of choice, as shown in figure 6.
@@
  • H-INT: Develop interfaces to suit the underlying interactions. Design an interface to best interact with the features of the tool, assuming that it provides good usability, there is no preference between step by step options, graph view windows and inspector editors.
@@
  • H-API: Integrate with application programming interfaces (APIs). The tool should be integrated as a part of an existing popular engine or editor, or at the very least should have full compatibility with said application. Technique interactions and processes must interact with each other while appropriately utilising the host game engine’s front-end and back-end APIs, as shown in figure 6.
Figure 8: Refined game asset design and development pipeline with heuristics resulting from research findings.